The Nutty Shoppe has 10 kg of mixed pecans and cashews, which sell for $ 8 per kilogram. Cashews alone sell for $8 per kilogram, and pecans sell for $9 per kilogram. How many kilograms of pecans are there in the mix?
How many kilograms of cashews are there in the mix?
Thanks

Answer by stanbon(75887)

The Nutty Shoppe has 10 kg of mixed pecans and cashews, which sell for $ 8 per kilogram. Cashews alone sell for $8 per kilogram, and pecans sell for $9 per kilogram. How many kilograms of pecans are there in the mix?
How many kilograms of cashews are there in the mix?
Let amount of cashews be "x" kg
Then amount of pecans is "10-x" kg
Value of the cashews is 8x dollars.
Value of the pecans is 9(10-x)=90-9x dollars
Value of mixture = 8(10)= 80 dollars
EQUATION:
value + value = value
8x+90-9x=80
-x =-10
x=10 kg (amount of cashews)
10-x=0 kg (amount of pecans}
Quite a mixture; don't you think.
Cheers,
Stan H.